OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x – 3y = –3 2x – 6y = –6

OpenStudy (anonymous):

1 0 –1 2

OpenStudy (ash2326):

@deesmoopth94 did you understand how to solve these type of questions?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

??

OpenStudy (ash2326):

Ok I'm gonna explain you we have \[5x – 3y = –3\] \[2x – 6y = –6\] Let's look at the coefficients of x and y We notice that y's coefficient in the two equations are -3 and -6 If we multiply the first by equation by 2 and subtract 2nd equation from it. We'd be able to eliminate y and can solve for x Multiply 1st by 2 we get \[10x-6y=-6\] now let's subtract 2nd equation from this we get \[10x-6y-(2x-6y)=-6-(-6)\] \[10x-6y-2x+6y=0\] \[10x-\cancel {6y}-2x+\cancel {6y}=0\] \[10x-2x=0\] we get \[8x=0\] or \[x= 0\] now to find y let's take any of the equations, say first \[5x – 3y = –3\] let's substitute x= 0 \[0-3y=-3\] divide both sides by -2 so we get y=1

OpenStudy (anonymous):

ohhhhh ok

OpenStudy (ash2326):

Did you understand?

OpenStudy (anonymous):

yes thank you